Exercise 5-17 Break-Even and Target Profit Analysis (L05-4, LO5-5, L05-6] Outback Outfitters sells recreational equipment. OnRequired 1 Required 2 Required 3 ] Required 4 At present, the company is selling 19,000 stoves per month. The sales manager iPlease put income statement in proper format with calculations explained. Thank you!
Answer #1

Answer

--Requirements asked: requirement 1 and 3

--Requirement 1

A Sale price $90
B Variable Cost $63
C = A - B Unit contribution margin $27
D Fixed Cost $132,300
E = D/C Break even point in units 4900 Answer
F = E x A Break even point in $ $441,000 Answer

--Requirement 3

Present Proposed
19000 stoves 23750 stoves
Total per unit Total per unit
Sale $1,710,000 $90 $1,923,750 $81
Variable cost $1,197,000 $63 $1,496,250 $63
Contribution margin $513,000 $27 $427,500 $18
Fixed cost $132,300 $132,300
Net Income or net operating income $380,700 $295,200

>Workings and calculations

Present Proposed
19000 stoves =19000+(19000*25%) stoves
Total per unit Total per unit
Sale =19000*90 90 =23750*81 =90-(90*10%)
Variable cost =19000*63 63 =23750*63 63
Contribution margin =1710000-1197000 =90-63 =1923750-1496250 =81-63
Fixed cost 132300 132300
Net Income or net operating income =513000-132300 =427500-132300